### v1

@@ Appendix A General Central Limit Theorem
The proof is identical to that of Theorem 4.3 in Gupta and Qu (2024) and is omitted. ∎
@@ Appendix B Proofs of Theorems
Denote $r_{\delta}=(r_{1\delta},\cdots,r_{n\delta})^{\prime}$. Note that $\hat{\xi}=\xi+(L^{\prime}P_{K_{1}}L)^{-1}L^{\prime}P_{K_{1}}u=(L^{\prime}P_{K_{1}}L)^{-1}L^{\prime}P_{K_{1}}(r_{\delta}+\epsilon)$ and $\hat{\alpha}=\alpha+R(L^{\prime}P_{K_{1}}L)^{-1}L^{\prime}P_{K_{1}}(r_{\delta}+\epsilon)$. Under $H_{0}:\alpha=0$, we have $\hat{\alpha}=R(L^{\prime}P_{K_{1}}L)^{-1}L^{\prime}P_{K_{1}}(r_{\delta}+\epsilon)$. Then we are going to investigate
@@
where $A_{11}=r_{\delta}^{\prime}(\widehat{\mathscr{M}}_{2}-\mathscr{M}_{2})r_{\delta}$, $A_{12}=r_{\delta}^{\prime}\mathscr{M}_{2}r_{\delta}$, $A_{21}=r_{\delta}^{\prime}(\widehat{\mathcal{M}}_{2}-\mathcal{M}_{2})v$, $A_{22}=r_{\delta}^{\prime}\mathcal{M}_{2}v$, ^$\widehat{\mathcal{M}}_{2}=\frac{1}{n}B^{\prime}K_{1}\widehat{\mathcal{V}}_{2}K_{1}^{\prime}B$ with $\widehat{\mathcal{V}}_{2}=(K_{1}^{\prime}K_{1})^{-1}K_{1}^{\prime}LT_{1}^{-1}R^{\prime}\widehat{\mathscr{D}}_{2}^{-1}RT_{1}^{-1}L^{\prime}K_{1}(K_{1}^{\prime}K_{1})^{-1}$, $\widehat{\mathscr{D}}_{2}=\frac{1}{n}R\widehat{\mathcal{U}}_{1}R^{\prime}$, $\widehat{\mathcal{U}}_{1}=T_{1}^{-1}L^{\prime}K_{1}(K_{1}^{\prime}K_{1})^{-1}\hat{\Xi}_{1}(K_{1}^{\prime}K_{1})^{-1}K_{1}^{\prime}LT_{1}^{-1}$, and ^$\widehat{\mathscr{M}}_{2}=\frac{1}{n}K_{1}\widehat{\mathcal{V}}_{2}K_{1}^{\prime}$. The remainder of the proof is similar to the proof of Theorem 5.1. ∎
@@
where $A_{11}=r_{G\delta}^{\prime}(\widehat{\mathscr{M}}_{3}-\mathscr{M}_{3})r_{G\delta},A_{12}=r_{G\delta}^{\prime}\mathscr{M}_{3}r_{G\delta}$, $A_{21}=r_{G\delta}^{\prime}(\widehat{\mathcal{M}}_{3}-\mathcal{M}_{3})v$, $A_{22}=r_{G\delta}^{\prime}\mathcal{M}_{3}v$, ^$\widehat{\mathcal{M}}_{3}=\frac{1}{n}B^{\prime}K_{2}\widehat{\mathcal{V}}_{3}K_{2}^{\prime}B$ with $\widehat{\mathcal{V}}_{3}=(K_{2}^{\prime}K_{2})^{-1}K_{2}^{\prime}FT_{2}^{-1}R^{\prime}\widehat{\mathscr{D}}_{3}^{-1}RT_{2}^{-1}F^{\prime}K_{2}(K_{2}^{\prime}K_{2})^{-1}$, $\widehat{\mathscr{D}}_{3}=\frac{1}{n}R\widehat{\mathcal{U}}_{2}R^{\prime}$, $\widehat{\mathcal{U}}_{2}=T_{2}^{-1}F^{\prime}K_{2}(K_{2}^{\prime}K_{2})^{-1}\allowbreak\hat{\Xi}_{2}(K_{2}^{\prime}K_{2})^{-1}K_{2}^{\prime}FT_{2}^{-1}$, and ^$\widehat{\mathscr{M}}_{3}=\frac{1}{n}K_{2}\widehat{\mathcal{V}}_{3}K_{2}^{\prime}$. The remainder of the proof is similar to the proof of Theorem 5.1. ∎
@@
because $\mathrm{E}(r_{i\lambda}^{2})=o(n^{-1})$ and $\mathrm{E}(r_{i\delta}^{2})=o(n^{-1})$, by Assumptions 16 and 17.
Next,
@@
Now, observe that
$\hat{\Xi}_{r,s}-\Xi_{r,s}=a_{1,rs}+a_{2,rs}+a_{3,rs},$ (B.8)
where
@@
where recall that $\sigma_{ij}$ is a typical element of $\Sigma$.
We will establish $a_{l,rs}=o_{p}(1),l=1,2,3$. To begin, write $r_{i}=r_{i\lambda}+r_{i\delta}$, $g_{i}^{\prime}$ for the $i$-th row of $G$ and note that
@@
Now, by Assumptions 15 and 16, $\left|r_{i}\right|=o_{p}(1/\sqrt{n})$, $\left\|g_{i}\right\|=O_{p}(\sqrt{d_{\gamma}})$ and, as in the proof of Theorem 2 of Kelejian and Prucha (2007), $\sum_{j=1}^{n}\left(1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})\right)\leq\ell$, all uniformly in $i$.
Then, we can write $\left|a_{1,rs}\right|\leq\sum_{l=1}^{7}a_{1l,rs}$, where, also using (B.7),
$a_{11,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}\left(1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})\right)\left\|g_{i}\right\|\left\|g_{j}\right\|\left\|\hat{\gamma}-\gamma\right\|^{2}\leq C\ell\cdot O_{p}\left(\frac{d_{\gamma}^{2}}{n}\right),$
$a_{12,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}\left(1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})\right)\left\|g_{i}\right\|\left\|\hat{\gamma}-\gamma\right\|\left|r_{j}\right|\leq C\ell\cdot O_{p}\left(\frac{d_{\gamma}}{n}\right),$
$a_{13,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}\left(1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})\right)\left|r_{i}\right|\left|r_{j}\right|\leq C\ell\cdot O_{p}\left(\frac{1}{n}\right),$
$a_{14,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}\left(1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})\right)\left\|g_{j}\right\|\left\|\hat{\gamma}-\gamma\right\|\left|\epsilon_{i}\right|\leq C\ell^{1-1/q}\cdot O_{p}\left(\frac{d_{\gamma}}{n^{1/2-1/q}}\right),$
$a_{15,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}\left(1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})\right)\left\|g_{i}\right\|\left\|\hat{\gamma}-\gamma\right\|\left|\epsilon_{j}\right|\leq C\ell^{1-1/q}\cdot O_{p}\left(\frac{d_{\gamma}}{n^{1/2-1/q}}\right),$
$a_{16,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}\left(1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})\right)\left|\epsilon_{i}\right|\left|r_{j}\right|\leq C\ell^{1-1/q}\cdot O_{p}\left(\frac{1}{n^{1/2-1/q}}\right),$
$a_{17,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}\left(1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})\right)\left|\epsilon_{j}\right|\left|r_{i}\right|\leq C\ell^{1-1/q}\cdot O_{p}\left(\frac{1}{n^{1/2-1/q}}\right),$
where the Hölder inequality is used to bound $a_{1j,rs},j\geq 4$, as in the proof of Theorem 2 of Kelejian and Prucha (2007). Therefore, we need $\frac{\ell d_{\gamma}^{2}}{n}=o_{p}(1)$ and $\frac{\ell^{1-1/q}d_{\gamma}}{n^{1/2-1/q}}=o_{p}(1)$ as $n\rightarrow\infty$. By Assumption 9(a), we have $\frac{\ell d_{\gamma}^{2}}{n}=o_{p}\left(n^{\eta-1}d_{\gamma}^{2}\right)$ and $\frac{\ell^{1-1/q}d_{\gamma}}{n^{1/2-1/q}}=o_{p}\left(n^{\eta(q-1)/q-(q-2)/2q}d_{\gamma}\right)$. We then have $a_{1,rs}=o_{p}(1)$ under the sufficient rate in (5.7) as $n\rightarrow\infty$.
$a_{2,rs}=o_{p}(1)$ and $a_{3,rs}=o_{p}(1)$ are established exactly as the negligibility of $b_{rs,n}$ and $c_{rs,n}$ is established in the proof of Theorem 1 of Kelejian and Prucha (2007). This concludes the proof. ∎

### v2

@@ Appendix A General Central Limit Theorem
${\left(v^{\prime}\mathscr{A}v-J\right)}/{\sqrt{2J}}=\sum_{s=1}^{n}\rho_{s}$, where
$\rho_{s}=L^{-1}a_{ss}\left(v_{s}^{2}-1\right)+2L^{-1}\mathbf{1}(s\geq 2)v_{s}\sum_{t<s}a_{st}v_{t},$
writing $L=\sqrt{2J}$. From Scott (1973), (A.2) follows if
$\sum_{s=1}^{n}\mathrm{E}\rho_{s}^{4}\overset{p}{\longrightarrow}0,\text{ as }n\rightarrow\infty,$ (A.3)
$\text{and }\sum_{s=1}^{n}\left[\mathrm{E}\left(\rho_{s}^{2}\left.{}\right|v_{t},t<s\right)-\mathrm{E}\left(\rho_{s}^{2}\right)\right]\overset{p}{\longrightarrow}0,\text{ as }n\rightarrow\infty.$ (A.4)
We show (A.3) first. Evaluating the LHS and using Assumption A yields
$\mathrm{E}\rho_{s}^{4}$ $\leq$ $CL^{-4}a_{ss}^{4}+CL^{-4}\sum_{t<s}a_{st}^{4}\leq CL^{-4}\left(\sum_{t\leq s}a_{st}^{2}\right)^{2}\leq CL^{-4}\left(b_{s}^{\prime}\mathscr{S}\sum_{t\leq s}b_{t}b_{t}^{\prime}\mathscr{S}b_{s}\right)^{2}$
$\leq$ $CL^{-4}\left(b_{s}^{\prime}\mathscr{S}^{2}b_{s}\right)^{2}=CL^{-4}\sum_{i,j,k=1}^{n}b_{is}b_{ks}s_{ij}s_{kj}$
$\leq$ $CL^{-4}\sum_{i,k=1}^{n}\left|b_{is}\right|\left|b_{ks}\right|\sum_{j=1}^{n}\left(s^{2}_{ij}+s^{2}_{kj}\right)=O_{p}\left(L^{-4}Jn^{-1}\left(\sum_{i=1}^{n}\left|b_{is}\right|\right)^{2}\right),$
so $\sum_{s=1}^{n}\mathrm{E}\rho_{s}^{4}=O_{p}\left(L^{-4}Jn^{-1}\sum_{s=1}^{n}\left(\sum_{i=1}^{n}\left|b_{is}\right|\right)^{2}\right)=O_{p}\left(L^{-4}J\right)$. Thus (A.3) is established. Note that $\mathrm{E}\left(\left.\rho_{s}^{2}\right|v_{t},t<s\right)=4L^{-2}\left\{\left(\mu_{4}-1\right)a_{ss}^{2}+2\mu_{3}\mathbf{1}(s\geq 2)\sum_{t<s}a_{st}a_{ss}v_{t}\right\}+4L^{-2}\mathbf{1}(s\geq 2)\left(\sum_{t<s}a_{st}v_{t}\right)^{2},$ where $\mathrm{E}v_{t}^{j}=\mu_{j},j=3,4$, and $\mathrm{E}\rho_{s}^{2}=4L^{-2}\left(\mu_{4}-1\right)a_{ss}^{2}+4L^{-2}\mathbf{1}(s\geq 2)\sum_{t<s}a_{st}^{2},$ so that (A.4) is bounded by
$CL^{-2}\sum_{s=2}^{n}\sum_{t<s}a_{st}a_{ss}v_{t}+C\left\{\sum_{s=2}^{n}\left(\sum_{t<s}a_{st}v_{t}\right)^{2}-\sum_{t<s}a_{st}^{2}\right\}.$ (A.5)
By transforming the range of summation, the square of the first term in (A.5) has expectation bounded by
$CL^{-4}\mathrm{E}\left(\sum_{t=1}^{n-1}\sum_{s=t+1}^{n}a_{st}a_{ss}v_{t}\right)^{2}\leq CL^{-4}\sum_{t=1}^{n-1}\left(\sum_{s=t+1}^{n}a_{st}a_{ss}\right)^{2},$ (A.6)
where the factor in parentheses on the RHS of (A.6) is
$\sum_{s,r=t+1}^{n}b_{s}^{\prime}\mathscr{S}b_{s}b_{s}^{\prime}\mathscr{S}b_{t}b_{r}^{\prime}\mathscr{S}b_{r}b_{r}^{\prime}\mathscr{S}b_{t}\leq\sum_{s,r=t+1}^{n}\left|b_{s}^{\prime}\mathscr{S}b_{s}b_{r}^{\prime}\mathscr{S}b_{r}\right|\left|b_{s}^{\prime}\mathscr{S}b_{t}\right|\left|b_{r}^{\prime}\mathscr{S}b_{t}\right|$
$\leq$ $C\sum_{s,r=t+1}^{n}\sum_{i,j,k,l=1}^{n}\left|b_{is}\right|\left|s_{ij}\right|\left|b_{jr}\right|\left|b_{ks}\right|\left|s_{lk}\right|\left|b_{kr}\right|\left|b_{s}^{\prime}\mathscr{S}b_{t}\right|\left|b_{r}^{\prime}\mathscr{S}b_{t}\right|$
$\leq$ $C\left(\sup_{i,j}\left|s_{ij}\right|\right)^{2}\left(\sup_{s\geq 1}\sum_{i=1}^{n}\left|b_{is}\right|\right)^{4}\sum_{s,r=t+1}^{n}\left|b_{s}^{\prime}\mathscr{S}b_{t}\right|\left|b_{r}^{\prime}\mathscr{S}b_{t}\right|$
$=$ $O_{p}\left(J^{2}n^{-2}\left(\sum_{s=t+1}^{n}\left|b_{t}^{\prime}\mathscr{S}b_{s}\right|\right)^{2}\right)=O_{p}\left(J^{2}n^{-2}\left(\sum_{s=t+1}^{n}\sum_{i,j=1}^{n}\left|b_{it}\right|\left|s_{ij}\right|\left|b_{js}\right|\right)^{2}\right),$
where we used Assumption A. Now, by Assumption A, $\sum_{s=t+1}^{n}\sum_{i,j=1}^{n}\left|b_{it}\right|\left|s_{ij}\right|\left|b_{js}\right|=O_{p}\left(\sup_{i,j}\left|s_{ij}\right|\sup_{t}\sum_{i=1}^{n}\left|b_{it}\right|\sum_{j=1}^{n}\sum_{s=t+1}^{n}\left|b_{js}\right|\right)=O_{p}\left(J\sup_{t}\sum_{i=1}^{n}\left|b_{it}\right|\right),$ and thus (A.6) is $O_{p}\left(L^{-4}J^{4}n^{-2}\sup_{t}\left(\sum_{i=1}^{n}\left|b_{it}\right|\right)\left(\sum_{i=1}^{n}\left(\sum_{t=1}^{n-1}\left|b_{it}\right|\right)\right)\right)=O_{p}\left(L^{-4}J^{4}n^{-1}\right)$, whence the first term in (A.5) is $O_{p}\left(J^{2}n^{-1}\right)$, which is negligible.
Again transforming the range of summation and using the elementary inequality $|a+b|^{2}\leq C\left(a^{2}+b^{2}\right)$, we can bound the square of the second term in (A.5) by
$C\left(\sum_{t=1}^{n-1}\sum_{s=t+1}^{n}a_{st}^{2}\left(v_{t}^{2}-1\right)\right)^{2}+C\left(2\sum_{t=1}^{n-1}\sum_{r=1}^{t-1}\sum_{s=t+1}^{n}a_{st}a_{sr}v_{t}v_{r}\right)^{2}.$ (A.7)
The expectations of the two terms in (A.7) are bounded by a constant times $e_{1}$ and $e_{2}$, respectively, where $e_{1}=\sum_{t=1}^{n-1}\left(\sum_{s=t+1}^{n}a_{st}^{2}\right)^{2},e_{2}=\sum_{t=1}^{n-1}\sum_{r=1}^{t-1}\left(\sum_{s=t+1}^{n}a_{st}a_{sr}\right)^{2}.$ Thus (A.7) is $O_{p}\left(e_{1}+e_{2}\right)$. Using Assumption A and elementary inequalities, $e_{2}$ is bounded by
$\sum_{t=1}^{n-1}\sum_{r=1}^{t-1}\sum_{s=t+1}^{n}\sum_{u=t+1}^{n}b_{s}^{\prime}\mathscr{S}b_{t}b_{s}^{\prime}\mathscr{S}b_{r}b_{u}^{\prime}\mathscr{S}b_{t}b_{u}^{\prime}\mathscr{S}b_{r}$
$=$ $O_{p}\left(L^{-4}\sum_{s,r,t,u=1}^{n}\sum_{i,j=1}^{n}\left|b_{ir}\right|\left|s_{ij}\right|\left|b_{js}\right|\sum_{i,j=1}^{n}\left|b_{ir}\right|\left|s_{ij}\right|\left|b_{ju}\right|\sum_{i,j=1}^{n}\left|b_{it}\right|\left|s_{ij}\right|\left|b_{js}\right|\right$.
$\times\left.\sum_{i,j=1}^{n}\left|b_{it}\right|\left|s_{ij}\right|\left|b_{ju}\right|\right)$
$=$ $O_{p}\left(L^{-4}pn^{-1}\sum_{s,r,t=1}^{n}\left(\sum_{i,j=1}^{n}\left|b_{ir}\right|\left|s_{ij}\right|\left|b_{js}\right|\right)\left(\sum_{i,j=1}^{n}\left|b_{ir}\right|\left|s_{ij}\right|\sum_{u=1}^{n}\left|b_{ju}\right|\right)\right$.
$\times$ $\left.\sum_{i,j=1}^{n}\left|b_{it}\right|\left|s_{ij}\right|\left|b_{js}\right|\sum_{i=1}^{n}\left|b_{it}\right|\sup_{u}\sum_{j=1}^{n}\left|b_{ju}\right|\right)$
$=$ $O_{p}\left(L^{-4}J^{2}n^{-2}\sum_{s,r=1}^{n}\left(\sum_{i,j=1}^{n}\left|b_{ir}\right|\left|s_{ij}\right|\left|b_{js}\right|\right)\sum_{i=1}^{n}\left|b_{ir}\right|\sum_{j=1}^{n}\left(\sum_{u=1}^{n}\left|b_{ju}\right|\right)\right$.
$\times$ $\left.\left(\sum_{i,j=1}^{n}\sum_{t=1}^{n}\left|b_{it}\right|\left|s_{ij}\right|\left|b_{js}\right|\right)\right)$
$=$ $O_{p}\left(L^{-4}J^{2}n^{-1}\sum_{i,j=1}^{n}\left(\sum_{r=1}^{n}\left|b_{ir}\right|\right)\left|s_{ij}\right|\left(\sum_{s=1}^{n}\left|b_{js}\right|\right)\left(\sup_{j}\sum_{i=1}^{n}\left|s_{ij}\right|\right)\sum_{j=1}^{n}\left|b_{js}\right|\right)$
$=$ $O_{p}\left(L^{-4}J^{2}n^{-1}\sup_{k}\sum_{i,j=1}^{n}\left|s_{ij}\right|\sum_{i=1}^{n}\left|s_{ik}\right|\right)=O_{p}\left(L^{-4}J^{2}n^{-1}\sup_{k}\sum_{i,j,\ell=1}^{n}\left|s_{ij}\right|\left|s_{\ell k}\right|\right)$
$=$ $O_{p}\left(L^{-4}J^{2}n^{-1}\sup_{k}\sum_{i,j,\ell=1}^{n}\left(s_{ij}^{2}+s_{\ell k}^{2}\right)\right)=O_{p}\left(L^{-4}J^{2}n^{-1}\sum_{i,j,\ell=1}^{n}\left(s_{ij}^{2}+s_{\ell j}^{2}\right)\right)$
$=$ $O_{p}\left(L^{-4}J^{2}n^{-1}\sum_{i,j=1}^{n}s_{ij}^{2}\right)=O_{p}\left(L^{-4}J^{2}\sup_{j}\sum_{i=1}^{n}s_{ij}^{2}\right)=O_{p}\left(Jn^{-1}\right),$
Similary $e_{1}$ is
$O_{p}\left(L^{-4}\sum_{t=1}^{n-1}\left\{\sum_{s=t+1}^{n}\left(\sum_{i,j=1}^{n}\left|s_{ij}\right|\left|b_{jt}\right|\left|b_{is}\right|\right)^{2}\right\}^{2}\right)$
$=$ $O_{p}\left(L^{-4}\left(\sup_{i,j}\left|s_{ij}\right|\right)^{4}\sum_{t=1}^{n-1}\left\{\sum_{s=t+1}^{n}\left(\sum_{i=1}^{n}\left|b_{is}\right|\sum_{j=1}^{n}\left|b_{jt}\right|\right)^{2}\right\}^{2}\right)$
$=$ $O_{p}\left((LJN)^{-4}\sum_{t=1}^{n-1}\left\{\sum_{s=t+1}^{n}\left(\sum_{i=1}^{n}\left|b_{is}\right|\right)^{2}\left(\sum_{j=1}^{n}\left|b_{jt}\right|\right)^{2}\right\}^{2}\right)$
$=$ $O_{p}\left((LJN)^{-4}\sum_{t=1}^{n-1}\left(\sum_{s=t+1}^{n}\left(\sum_{i=1}^{n}\left|b_{is}\right|\right)^{2}\right)^{2}\left(\sum_{j=1}^{n}\left|b_{jt}\right|\right)^{4}\right)$
$=$ $O_{p}\left((LJN)^{-4}\left(\sum_{t=1}^{n-1}\sum_{j=1}^{n}\left|b_{jt}\right|\right)\left(\sum_{s=t+1}^{n}\sum_{i=1}^{n}\left|b_{is}\right|\right)^{2}\right$.
$\times\left.\sup_{s}\left(\sum_{i=1}^{n}\left|b_{is}\right|\right)^{2}\sup_{t}\left(\sum_{j=1}^{n}\left|b_{jt}\right|\right)^{3}\right)$
$=$ $O_{p}\left((L^{-4}J^{4}n^{-1}\right)=O_{p}\left(J^{2}n^{-1}\right).$
@@ Appendix B Proofs of Theorems
Denote $r_{\delta}=(r_{1\delta},\cdots,r_{n\delta})^{\prime}$. Note that $\hat{\xi}=\xi+(L^{\prime}P_{K_{1}}L)^{-1}L^{\prime}\\ P_{K_{1}}u=(L^{\prime}P_{K_{1}}L)^{-1}L^{\prime}P_{K_{1}}(r_{\delta}+\epsilon)$ and $\hat{\alpha}=\alpha+R(L^{\prime}P_{K_{1}}L)^{-1}L^{\prime}P_{K_{1}}(r_{\delta}+\epsilon)$. Under $H_{0}:\alpha=0$, we have $\hat{\alpha}=R(L^{\prime}P_{K_{1}}L)^{-1}L^{\prime}P_{K_{1}}(r_{\delta}+\epsilon)$. Then we are going to investigate
@@
where $A_{11}=r_{\delta}^{\prime}(\widehat{\mathscr{M}}_{2}-\mathscr{M}_{2})r_{\delta}$, $A_{12}=r_{\delta}^{\prime}\mathscr{M}_{2}r_{\delta}$, $A_{21}=r_{\delta}^{\prime}(\widehat{\mathcal{M}}_{2}-\mathcal{M}_{2})v$, $A_{22}=r_{\delta}^{\prime}\mathcal{M}_{2}v$, $\widehat{\mathcal{M}}_{2}=\frac{1}{n}B^{\prime}K_{1}\widehat{\mathcal{V}}_{2}K_{1}^{\prime}B$ with $\widehat{\mathcal{V}}_{2}=(K_{1}^{\prime}K_{1})^{-1}K_{1}^{\prime}LT_{1}^{-1}R^{\prime}\widehat{\mathscr{D}}_{2}^{-1}RT_{1}^{-1}L^{\prime}K_{1}(K_{1}^{\prime}K_{1})^{-1}$, $\widehat{\mathscr{D}}_{2}=\frac{1}{n}R\widehat{\mathcal{U}}_{1}R^{\prime}$, $\widehat{\mathcal{U}}_{1}=T_{1}^{-1}L^{\prime}K_{1}(K_{1}^{\prime}K_{1})^{-1}\hat{\Xi}_{1}(K_{1}^{\prime}K_{1})^{-1}K_{1}^{\prime}LT_{1}^{-1}$, and $\widehat{\mathscr{M}}_{2}=\frac{1}{n}K_{1}\widehat{\mathcal{V}}_{2}K_{1}^{\prime}$. The remainder of the proof is similar to the proof of Theorem 5.1. ∎
@@
where $A_{11}=r_{G\delta}^{\prime}(\widehat{\mathscr{M}}_{3}-\mathscr{M}_{3})r_{G\delta},A_{12}=r_{G\delta}^{\prime}\mathscr{M}_{3}r_{G\delta}$, $A_{21}=r_{G\delta}^{\prime}(\widehat{\mathcal{M}}_{3}-\mathcal{M}_{3})v$, $A_{22}=r_{G\delta}^{\prime}\mathcal{M}_{3}v$, $\widehat{\mathcal{M}}_{3}=\frac{1}{n}B^{\prime}K_{2}\widehat{\mathcal{V}}_{3}K_{2}^{\prime}B$, $\widehat{\mathcal{V}}_{3}=(K_{2}^{\prime}K_{2})^{-1}K_{2}^{\prime}FT_{2}^{-1}R^{\prime}\widehat{\mathscr{D}}_{3}^{-1}RT_{2}^{-1}F^{\prime}K_{2}(K_{2}^{\prime}K_{2})^{-1}$, $\widehat{\mathscr{D}}_{3}=\frac{1}{n}R\widehat{\mathcal{U}}_{2}R^{\prime}$, $\widehat{\mathcal{U}}_{2}=T_{2}^{-1}F^{\prime}K_{2}(K_{2}^{\prime}K_{2})^{-1}\allowbreak\hat{\Xi}_{2}(K_{2}^{\prime}K_{2})^{-1}K_{2}^{\prime}FT_{2}^{-1}$, and $\widehat{\mathscr{M}}_{3}=\frac{1}{n}K_{2}\widehat{\mathcal{V}}_{3}K_{2}^{\prime}$. The remainder of the proof is similar to the proof of Theorem 5.1. ∎
@@
because $\mathrm{E}(r_{i\lambda}^{2})=o(n^{-1})$ and $\mathrm{E}(r_{i\delta}^{2})=o(n^{-1})$, by Assumptions 16 and 17. Next,
@@
Now, observe that $\hat{\Xi}_{r,s}-\Xi_{r,s}=a_{1,rs}+a_{2,rs}+a_{3,rs},$ where
@@
where recall that $\sigma_{ij}$ is a typical element of $\Sigma$. We will show $a_{l,rs}=o_{p}(1),l=1,2,3$. To begin, write $r_{i}=r_{i\lambda}+r_{i\delta}$, $g_{i}^{\prime}$ for the $i$-th row of $G$ and note that
@@
Now, by Assumptions 15 and 16, $\left|r_{i}\right|=o_{p}(1/\sqrt{n})$, $\left\|g_{i}\right\|=O_{p}(\sqrt{d_{\gamma}})$ and, as in the proof of Theorem 2 of Kelejian and Prucha (2007), $\sum_{j=1}^{n}\left(1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})\right)\leq\ell$, all uniformly in $i$. Then, we can write $\left|a_{1,rs}\right|\leq\sum_{l=1}^{7}a_{1l,rs}$, where, writing $\pi_{M}=1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})$ and also using (B.7),
$a_{11,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}\pi_{M}\left\|g_{i}\right\|\left\|g_{j}\right\|\left\|\hat{\gamma}-\gamma\right\|^{2}\leq C\ell\cdot O_{p}\left({d_{\gamma}^{2}}/{n}\right),$
$a_{12,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}\pi_{M}\left\|g_{i}\right\|\left\|\hat{\gamma}-\gamma\right\|\left|r_{j}\right|\leq C\ell\cdot O_{p}\left({d_{\gamma}}/{n}\right),$
$a_{13,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}\pi_{M}\left|r_{i}\right|\left|r_{j}\right|\leq C\ell\cdot O_{p}\left({1}/{n}\right),$
$a_{14,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}\pi_{M}\left\|g_{j}\right\|\left\|\hat{\gamma}-\gamma\right\|\left|\epsilon_{i}\right|\leq C\ell^{1-1/q}\cdot O_{p}\left({d_{\gamma}}/{n^{1/2-1/q}}\right),$
$a_{15,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}\pi_{M}\left\|g_{i}\right\|\left\|\hat{\gamma}-\gamma\right\|\left|\epsilon_{j}\right|\leq C\ell^{1-1/q}\cdot O_{p}\left({d_{\gamma}}/{n^{1/2-1/q}}\right),$
$a_{16,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}\pi_{M}\left|\epsilon_{i}\right|\left|r_{j}\right|\leq C\ell^{1-1/q}\cdot O_{p}\left({1}/{n^{1/2-1/q}}\right),$
$a_{17,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}\pi_{M}\left|\epsilon_{j}\right|\left|r_{i}\right|\leq C\ell^{1-1/q}\cdot O_{p}\left({1}/{n^{1/2-1/q}}\right),$
where the Hölder inequality is used to bound $a_{1j,rs},j\geq 4$, as in the proof of Theorem 2 of Kelejian and Prucha (2007). Therefore, we need $\frac{\ell d_{\gamma}^{2}}{n}=o_{p}(1)$ and $\frac{\ell^{1-1/q}d_{\gamma}}{n^{1/2-1/q}}=o_{p}(1)$ as $n\rightarrow\infty$. By Assumption 9(a), we have $\frac{\ell d_{\gamma}^{2}}{n}=o_{p}\left(n^{\eta-1}d_{\gamma}^{2}\right)$ and $\frac{\ell^{1-1/q}d_{\gamma}}{n^{1/2-1/q}}=o_{p}\left(n^{\eta(q-1)/q-(q-2)/2q}d_{\gamma}\right)$. We then have $a_{1,rs}=o_{p}(1)$ under the sufficient rate in (5.7) as $n\rightarrow\infty$. $a_{2,rs}=o_{p}(1)$ and $a_{3,rs}=o_{p}(1)$ are established exactly as the negligibility of $b_{rs,n}$ and $c_{rs,n}$ is established in the proof of Theorem 1 of Kelejian and Prucha (2007). This concludes the proof. ∎